OpenStudy (anonymous):

The population of a small town, P, as a function of time, t, in years past 1940 is given below. P = 2,369 + 125t For which of the following years was the population of the town 5,369? 1964 1924 1934 1974

OpenStudy (anonymous):

@misty1212

OpenStudy (misty1212):

\[P = 2,369 + 125t =5369\]

OpenStudy (misty1212):

\[t=24\] \[1940+24=1964\]